OpenStudy (anonymous):

How do you change powers with negative exponents to powers with positive exponents?

OpenStudy (phi):

you "flip them" \[ x^{-a} = \frac{x^{-a}}{1}=\frac{1}{x^a} \] also \[ \frac{1}{x^{-a}}=\frac{x^a}{1} = x^a\]

OpenStudy (anonymous):

Thank You